ganeshie8 (ganeshie8):

how to evaluate this \[\large \sum\limits_{m,n=1}^3 (mn) = ?\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

1*(1+2+3) + 2*(1+2+3) + 3*(1+2+3) ?

OpenStudy (anonymous):

\[= \left( \sum\limits_{n=1}^3 n\right)\left( \sum\limits_{m=1}^3 m\right) \]

ganeshie8 (ganeshie8):

Nice :) the one i wrote is a short form for that product is it ?

OpenStudy (anonymous):

which is the same as @INewton wrote

ganeshie8 (ganeshie8):

Oh I see...

OpenStudy (dan815):